OpenStudy (anonymous):

on a street that is 75m long, a fire hydrant is placed to that it is 1m farther from the west end of the street than from the east end. how far is the fire hydrant from each side of the street?

OpenStudy (anonymous):

plzzzz help!!!!!

OpenStudy (anonymous):

Set up a system of equations with x being the distance from the west side and y being the distance from the east side: x + y = 75 x = y + 1 Solve the system using substitution: (y + 1) + y = 75 2y + 1 =75 2y = 74 y = 37 x = (37) + 1 x = 38 Now you know that distance from the west side is 38 meters and the distance from the east side is 38 meters.